OpenStudy (love_to_love_you):

Find the equation of the tangent line to the function f(x) = x 4 − x + 3 at a = 1.

OpenStudy (love_to_love_you):

y = 3x + 8 y = 3x + 6 y = 3x − 6 y = 3x